OpenStudy (anonymous):

how to find domain for this question y= sqrt(x^2+x-2)

OpenStudy (anonymous):

A domain is what x is allowed to be in your function You can't take the square root of a negative. So your domain is the solution to x^2 + x - 2 >= 0

OpenStudy (anonymous):

-2 and 1

OpenStudy (anonymous):

Those are the solutions to x^2 + x - 2 = 0 Remember we are solving x^2 + x - 2 >= 0 So you get (x + 2)(x - 1) >= 0 Now set up a number line to get your solution ranges. So the domain is x <= -2 or x >= 1

OpenStudy (anonymous):

ok...i know....thanks u... coz this is sqrt...so just take positive part only right....